OpenStudy (anonymous):

Could someone please help me understand this problem... How would you write 19t in a word phrase?

jimthompson5910 (jim_thompson5910):

19t means "19 times t" where t is some unknown number (it's a variable)

jimthompson5910 (jim_thompson5910):

so you could say "19t" could represent "the product of 19 and t" or "the product of 19 and some unknown number"

OpenStudy (anonymous):

Thank you so much!

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

I have another one im stuck on, do you mind helping?

jimthompson5910 (jim_thompson5910):

sure, go for it

OpenStudy (anonymous):

\[\frac{ y }{ 11 }-3\]

OpenStudy (anonymous):

wrote out into a word phrase

jimthompson5910 (jim_thompson5910):

y/11 means "y divided by 11"

jimthompson5910 (jim_thompson5910):

btw, \[\large \frac{y}{11} = y/11\]

jimthompson5910 (jim_thompson5910):

so y/11 - 3 means "y divided by 11 minus 3" and you could say something like "the quotient of y and 11 minus 3" or "3 less than the quotient of y and 11"

OpenStudy (anonymous):

so it wouldn't have been a fraction just division ?

jimthompson5910 (jim_thompson5910):

a fraction is division, they are one in the same really

jimthompson5910 (jim_thompson5910):

example 1/2 = 0.5 one half = 0.5 you divide 1 into 2 to get 0.5 (use long division or a calculator)

OpenStudy (anonymous):

i think i kind of understand, i usually get quite panicked when i start to do my school work and see fractions! ive never been good when it comes to fractions

jimthompson5910 (jim_thompson5910):

keep practicing with them and you'll get better